# Multilinear Wavelet Compact T(1) Theorem

### v1

@@ 5. Compactness of paraproducts
Let $\vartheta\in C^{\infty}_{0}(B(0,1))$ with $\int\vartheta(x)\,{\rm d}x=1$. For each $z\in Z^{d}$, set $\vartheta_{z}={\mathsf{Sy}}_{z}\vartheta$. Given $\mathfrak{b}\in\mathrm{BMO}(\mathbb{R}^{d})$, define the $(m+1)$-linear form $\Pi_{\mathfrak{b}}$ by
@@
$\Pi_{b}$ is called a paraproduct form with symbol $\mathfrak{b}$. Any $m$-linear operator in $\mathsf{T}_{\Pi_{b}}$ is called a paraproduct with symbol $\mathfrak{b}$, which we denote by $S_{\mathfrak{b}}$.
@@
Before proving this, let us review the standard boundedness theory of paraproducts, analogous to Proposition 2.8 for wavelet forms. To do so, it is convenient to view $\Pi_{\mathfrak{b}}$ as an $(m+2)$-linear form, where the extra input function is $\mathfrak{b}$ itself. In fact, in this way $\Pi_{\mathfrak{b}}(f_{0},\ldots,f_{m})=U(\mathfrak{b},f_{0},\ldots,f_{m})$ where $U$ is a canonical $(m+2)$-linear wavelet form with $\left|\varepsilon_{z}\right|\lesssim 1$. Such forms are cancellative in the first and second positions, for which a slight strengthening of (2.7) and (2.8) holds:
(5.1) $T\in\mathsf{T}_{V_{\mathfrak{b}}},\quad V_{\mathfrak{b}}(f_{0},\ldots,f_{m})=U(\mathfrak{b},f_{0},\ldots,f_{m}),$
$\left\|T\right\|_{\vec{p},r}\leq C_{\vec{p},r}\left\|\mathfrak{b}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad\left\|T\right\|_{\infty,\mathrm{BMO}}\leq C_{\infty}\left\|\mathfrak{b}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad(\vec{p},r)\in Q.$
@@
$\langle P_{M}\mathfrak{b},\phi_{z}\rangle=\sum_{I\in{\mathcal{D}}_{M}}|I|\langle\mathfrak{b},\psi_{I}\rangle\langle\psi_{I},\phi_{z}\rangle=\left\langle\mathfrak{b},\sum_{I\in{\mathcal{D}}_{M}}|I|\langle\psi_{I}^{1},\phi_{z}\rangle\psi_{I}^{1}\right\rangle.$
@@
$\tilde{\psi}_{\zeta}=\psi_{I},\quad H(\zeta,z)=\frac{|I|}{\mu(I\times(\tfrac{\ell(I)}{2},\ell(I)])}\langle\psi_{I}^{1},\phi_{z}\rangle,$
@@ 6. Main result and proof
Suppose $\Lambda$ is an $(m+1)$-linear CZ form, with associated symbols $\mathfrak{b}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$, $j=0,1,\ldots,m$ satisfying (1.2). The following are equivalent.
@@
  • B.
    There exist compact wavelet forms $\{U_{k}\}_{k=1}^{K_{m}}$ and compact paraproduct forms $\Pi_{\mathfrak{b}_{j}}$ such that for all $f_{j}\in{\mathcal{S}}(\mathbb{R}^{d})$,
@@
  • C.
    Each element of $\ \mathsf{T}_{\Lambda}$ is a compact CZO.
@@
To show A. implies B. we isolate the cancellative part of $\Lambda$, namely
@@
and since $\phi$ is cancellative, $\Pi^{*,j}_{\mathfrak{b}}(\varphi,1,\ldots,1)=0$ for $j=1,\ldots,m$. We want to apply Proposition 3.9 to $\Lambda_{\mathsf{c}}$ so we must establish that it is a compact CZ form. Since $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, each $S\in\mathsf{T}_{\Pi_{\Lambda}}$ is compact by Proposition 5.2, so by Proposition 3.6, $\Pi_{\Lambda}$ is a compact CZ form. Since we also know that $\Lambda$ is a compact CZ form, $\Lambda_{\mathsf{c}}$ must indeed be compact, and B. follows by applying Proposition 3.9 to $\Lambda_{\mathsf{c}}$. B. implies C. is a consequence of Propositions 4.1 and 5.2. Finally, C. implies A. is the content of Proposition 3.6. ∎
A fourth equivalent condition in Theorem 2 would be the compactness of the adjoint operators on weighted Lebesgue spaces for weights belonging to the multilinear Muckenhoupt classes [lerner-et-al, nieraeth2019quantitative]. If one is only concerned with weights corresponding to $p_{j}<\infty$, then this follows immediately from [COY]. Alternatively, to handle the full range of $Q$, including when some $p_{j}=\infty$, one can verify the conditions of the weighted Riesz-Kolmogorov theorem [COY]*Proposition 2.9 as we did in Propositions 4.1 and 5.2 by hand. Furthermore, in [COY], there are provided examples of multilinear operators which are compact on these weighted spaces.
We note that Theorem 2 applies to the linear case as well and a few simplifications can be made due to the greater symmetry enjoyed in this setting. Let us restate Theorem 2 when $m=1$ for additional clarity.
@@
  • B.
    There exists a compact wavelet form $U$ such that for all $f,g\in{\mathcal{S}}(\mathbb{R}^{d})$,
@@
  • D.
    $T(1),T^{*}(1)\in\mathrm{CMO}(\mathbb{R}^{d})$ and
    $\lim_{z\to\infty}t^{\frac{d}{2}}\left\|T\phi_{z}\right\|_{L^{2}(\mathbb{R}^{d})}=0.$
@@
To compare A. with its $m$-linear analogue, one only needs to note that $z\in\mathbb{B}_{M}(\zeta)$ is equivalent to $\zeta\in\mathbb{B}_{M^{\prime}}(z)$ as long as $M\sim M^{\prime}$. Therefore, A. implies B. follows from Theorem 2. B. implies C. is a special case of Theorem 2. C. implies D. since $t^{\frac{d}{2}}\phi_{z}\to 0$ weakly in $L^{2}(\mathbb{R}^{d})$ as $z\to\infty$. Finally, D. implies A. by Cauchy-Schwarz. ∎

### v2

@@ 5. Compactness of paraproducts
Let $\vartheta\in{\mathrm{C}}^{\infty}_{0}(B(0,1))$ with $\int\vartheta(x)\,{\rm d}x=1$. For each $z\in Z^{d}$, set $\vartheta_{z}={\mathsf{Sy}}_{z}\vartheta$. Given $\mathfrak{b}\in\mathrm{BMO}(\mathbb{R}^{d})$, define the $(m+1)$-linear form $\Pi_{\mathfrak{b}}$ by
@@
$\Pi_{b}$ is called a paraproduct form with symbol $\mathfrak{b}$. Any $m$-linear operator $S_{\mathfrak{b}}$ such that
$\left\langle f_{0},S_{\mathfrak{b}}(f_{1},\ldots,f_{m})\right\rangle=\Pi_{\mathfrak{b}}^{\sigma}(f_{0},\ldots,f_{m})$
for some $\sigma\in\mathrm{S}_{m+1}$ is called a paraproduct with symbol $\mathfrak{b}$.
@@
Our proof of Proposition 5.2 will argue along the same lines as our proof of Proposition 4.1 from the previous section. For different line of reasoning, see [BLOT24a].
To prepare for the proof, let us review the standard boundedness theory of paraproducts, analogous to Proposition 2.8 for wavelet forms. To do so, it is convenient to view $\Pi_{\mathfrak{b}}$ as an $(m+2)$-linear form, where the extra input function is $\mathfrak{b}$ itself. In fact, in this way $\Pi_{\mathfrak{b}}(f_{0},\ldots,f_{m})=U(\mathfrak{b},f_{0},\ldots,f_{m})$ where $U$ is a canonical $(m+2)$-linear wavelet form with $\left|\varepsilon_{z}\right|\lesssim 1$. Such forms are cancellative in the first and second positions, for which a slight strengthening of (2.9) and (2.10) holds: If $T$ satisfies
$\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle=U(\mathfrak{b},f_{\sigma(0)},\ldots,f_{\sigma(m)})$
for some $\sigma\in\mathrm{S}_{m+1}$ and canonical wavelet form $U$, then
(5.1) $\left\|T\right\|_{\vec{p},r}\leq C_{\vec{p},r}\left\|\mathfrak{b}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad\left\|T\right\|_{\infty,\mathrm{BMO}}\leq C_{\infty}\left\|\mathfrak{b}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad(\vec{p},r)\in Q.$
@@
$\langle P_{M}\mathfrak{b},\phi_{z}\rangle=\sum_{I\in{\mathcal{D}}_{M}}|I|\langle\mathfrak{b},\psi_{I}\rangle\langle\psi_{I},\phi_{z}\rangle=\left\langle\mathfrak{b},\sum_{I\in{\mathcal{D}}_{M}}|I|\langle\psi_{I},\phi_{z}\rangle\psi_{I}\right\rangle.$
@@
$\tilde{\psi}_{\zeta}=\psi_{I},\quad H(\zeta,z)=\frac{|I|}{\mu(I\times(\tfrac{\ell(I)}{2},\ell(I)])}\langle\psi_{I},\phi_{z}\rangle,$
@@ 6. Main result and proof
Suppose $T$ is an $m$-linear CZO, $\Lambda$ the associated $(m+1)$-linear form, with $\mathfrak{b}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$, $j=0,1,\ldots,m$ satisfying (1.2). The following are equivalent.
@@
    $T$ is a compact CZO in the reflexive range.
  • B.
    For any $\varphi^{1},\ldots,\varphi^{m}\in L^{\infty}$ with compact support, and every $j=0,1,\ldots,m$,
    $\lim_{z\to\infty}t^{md-\frac{d}{2}}\left\|T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})\right\|_{L^{2}(\mathbb{R}^{d})}=0,\quad\varphi_{z}^{i}={\mathsf{Sy}}_{z}\varphi^{i}.$
  • C.
@@
  • D.
    There exist compact wavelet forms $\{U_{k}\}_{k=1}^{K_{m}}$ and compact paraproduct forms $\{\Pi_{\mathfrak{b}_{j}}\}_{j=0}^{m}$ such that for all $f_{j}\in{\mathcal{S}}(\mathbb{R}^{d})$,
@@
  • E.
    $T^{*,j}$ is a compact CZO for each $j=0,1,\ldots,m$.
@@
It is trivial that E. implies A., and the path (A. $\implies$ B. $\implies$ C.) is precisely the content of Proposition 3.8. To show C. implies D. we isolate the cancellative part of $\Lambda$, namely
@@
and since $\phi$ is cancellative, $\Pi^{*,j}_{\mathfrak{b}}(\varphi,1,\ldots,1)=0$ for $j=1,\ldots,m$. We want to apply Proposition 3.11 to $\Lambda_{\mathsf{c}}$ so we must establish that it is a compact CZ form. Since $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, by Remarks 3.3 and 3.4, $\Pi_{\Lambda}$ is a compact CZ form. Since we also know that $\Lambda$ is a compact CZ form, $\Lambda_{\mathsf{c}}$ must indeed be compact, and B. follows by applying Proposition 3.11 to $\Lambda_{\mathsf{c}}$. D. implies E. is a consequence of Propositions 4.1 and 5.2.∎
A sixth equivalent condition in Theorem 2 would be the compactness of the adjoint operators on weighted Lebesgue spaces for weights belonging to the multilinear Muckenhoupt classes [lerner-et-al, nieraeth2019quantitative]. If one is only concerned with weights corresponding to $p_{j}<\infty$, then this follows immediately from [COY]. Alternatively, to handle the full range of $Q$, including when some $p_{j}=\infty$, one can verify the conditions of the weighted Riesz-Kolmogorov theorem [COY]*Proposition 2.9 as we did in Propositions 4.1 and 5.2 by hand.
Furthermore, let us briefly recall the notion of a compact CZ kernel, which has been ubiquitous in the compact CZO literature until [MS23] recently removed this condition. A compact CZ kernel satisfies the kernel estimates (3.1) and (3.2), but in addition to the uniform bound provided by $C_{K}$, one imposes an additional prefactor on the RHS,
$o\left(\max\left\{\left|x_{0}-\mathbf{x}\right|_{2},\left|x_{0}-\mathbf{x}\right|_{2}^{-1},\left|x_{0}+\mathbf{x}\right|_{2}^{-1}\right\}\right).$
By standard arguments [tao-notes]*Lecture 7, if one assumes $T$ has a compact CZ kernel and satisfies the diagonal weak compactness property
(6.1) $\lim_{z\to\infty}\Lambda(\mathbf{1}_{B(w,t)},\ldots,\mathbf{1}_{B(w,t)})=0,$
then condition B. in Theorem 2 holds from whence all the other conclusions follow. This perspective, using the compact kernel and (6.1), was recently undertaken in [CLSY24] in the bilinear setting.
We finally note that Theorem 2 applies to the linear case as well and a few simplifications can be made due to the greater symmetry enjoyed in this setting. Let us restate Theorem 2 when $m=1$ for additional clarity.
@@
    $T$ is compact on $L^{2}(\mathbb{R}^{d})$.
  • B.
    For any $\varphi\in L^{\infty}_{c}$,
    $\lim_{z\to\infty}t^{\frac{d}{2}}\left\|T(\varphi_{z})\right\|_{L^{2}}+t^{\frac{d}{2}}\left\|T^{*}(\varphi_{z})\right\|_{L^{2}}=0,\quad\varphi_{z}={\mathsf{Sy}}_{z}\varphi.$
  • C.
@@
  • D.
    There exists a compact wavelet form $U$ and compact paraproduct forms $\Pi_{T(1)},\Pi_{T^{*}(1)}$ such that for all $f,g\in{\mathcal{S}}(\mathbb{R}^{d})$,
@@
Only A. and C. differ from their multilinear statements in Theorem 2. To compare C. with its $m$-linear analogue, one only needs to note that $z\in\mathbb{B}_{M}(\zeta)$ is equivalent to $\zeta\in\mathbb{B}_{M^{\prime}}(z)$ as long as $M\sim M^{\prime}$. Therefore, (B. $\implies$ C. $\implies$ D. $\implies$ E.) follows from Theorem 2. E. trivially implies A., so it remains to establish A. implies B. where the true force of the linear situation is used. Namely, $t^{\frac{d}{2}}\phi_{z}$ converges weakly to zero and therefore, since $T$ is compact $TT^{*}(t^{\frac{d}{2}}\phi_{z})$ and $T^{*}T(t^{\frac{d}{2}}\phi_{z})$ both converge strongly to zero, which implies B. ∎
Finally, we conclude with a short, simple application.
Example
Given functions $a,b\in L^{\infty}(\mathbb{R}^{d})$, consider the operators of multiplication on the Fourier and space sides by
[…]
It is easy to see that unless $a$ or $b$ is identically zero, such operators can never be compact. This even persists for the multilinear analogues, in particular the $m$-linear Fourier multiplier
$T_{a}(f_{1},\ldots,f_{m})(x)=\int_{\mathbb{R}^{md}}a(\xi)\prod_{j=1}^{m}\hat{f}_{j}(\xi_{j})e^{ix(\xi_{1}+\cdots+\xi_{m})}\,d\xi,\quad\xi=(\xi_{1},\ldots,\xi_{m}),\xi_{j}\in\mathbb{R}^{d}.$
Nonetheless, by the uncertainty principle, we expect compositions of $T_{a}$ and $M_{b}$ to be compact. The most well-known example is in the linear case when $a,b\in L^{2}(\mathbb{R}^{d})$. Then trivially $T_{a}M_{b}$ is Hilbert-Schmidt under no smoothness assumptions on $a$ and $b$.
We must impose some conditions on $a$ and functions $b_{0},\ldots,b_{m}$ so that suitable compositions of $T_{a}$ and $M_{b_{j}}$ are CZOs. Assume there exists $\delta>0$ such that $b_{j}\in{\mathrm{C}}^{\delta}(\mathbb{R}^{d})$ and $a$ belongs to ${\mathrm{C}}^{\infty}(\mathbb{R}^{md}\setminus\{0\})$ such that
(6.2) $\left|\partial^{\alpha}a(\xi)\right|\leq C_{\alpha}|\xi|^{-|\alpha|},\quad\left|b_{j}(x)\right|+\frac{\left|b_{j}(x)-b_{j}(x+h)\right|}{\left|h\right|^{\delta}}\lesssim(1+\left|x\right|)^{-\delta}.$
It is a routine exercise to check that under these conditions, the operator
$S(f_{1},\ldots,f_{m})=M_{b_{0}}T_{a}(M_{b_{1}}f_{1},\ldots,M_{b_{m}}f_{m})$
is an $m$-linear CZO satisfying $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$. Therefore, by Theorem 2, compactness of $S$ is characterized by $\mathsf{W}_{\Lambda}(z)\to 0$. To check that, we examine quantities of the form
(6.3) $\int_{\mathbb{R}^{md}}a(\xi_{1},\ldots,\xi_{m})\widehat{c_{0}g^{0}_{z_{0}}}(\xi_{1}+\cdots+\xi_{m})\prod_{j=1}^{m}\widehat{c_{j}g^{j}_{z_{j}}}(\xi_{j})\,d\xi,\quad z_{j}\in\mathbb{B}(z),$
where $(c_{0},\ldots,c_{m})$ is some permutation of $(b_{0},\ldots,b_{m})$ and each $g^{i}$ is either $\phi$, $\psi^{1}$, or $\psi^{3}$. One can rewrite, for any $\eta\in\mathbb{R}^{d}$,
$\widehat{c_{j}g^{j}_{z_{j}}}(\eta)=\left\langle e^{i\eta\cdot}c_{j},g^{j}_{z_{j}}\right\rangle$
to see that for each $\xi$, the integrand in (6.3) goes to zero as $z\to\infty$. Then, if $a\in L^{1}(\mathbb{R}^{md})$, by dominated convergence, the entire quantity (6.3) goes to zero, and $S$ is a compact CZO.